# Revisiting the Identification of the Conduct Parameter in Homogeneous Goods Markets

### v2

@@ 3.2 Main Result
We are going to show the following theorem, which is the formal version of the result in the introduction:
@@
Given Assumption 1, 2, 3, and 4, the conduct parameter $\theta$ cannot be identified from data on price, quantity, and other exogenous variables alone if and only if the industry inverse demand function is given as
@@
where $\alpha\neq-\frac{1}{\theta}$ is a constant, and $r:\mathbb{R}^{K^{d}}\rightarrow\mathbb{R}$ and $s:\mathbb{R}_{+}\rightarrow\mathbb{R}$ are twice continuously differentiable functions
@@
As in Lau’s claim, the demand shifters affect the inverse demand function only through a function $r$. Therefore, when the demand shifter is a vector, the inverse demand function is separable. However, our result tells more about the non-identification holds under what type of separable function and has a clear economic interpretation. Again, our result has a special case where the identification always holds ($\alpha=-\frac{1}{\theta}$). As we will discuss, the identification is always achieved because any demand shifter works as a demand rotation instrument in this case. Except in this case, observe that the demand shifter changes only the slope of the inverse demand function when $\alpha\neq 0$ or changes only the intercept of the inverse demand function when $\alpha=0$. Recall that a demand rotation instrument can change the intercept and the slope of the inverse demand function simultaneously without changing the equilibrium quantity. Therefore, we cannot have any demand rotation instrument under (15). The result implies that the identification holds if and only if the inverse demand function has demand shifters that work as demand rotation instruments. This clearly reinforces the idea of bresnahanOligopoly1982 using demand rotation instruments to identify the conduct parameter in more general settings. It also reinforces the sufficiency of demand rotation instruments for the identification in the model with a linear inverse demand and a linear marginal cost as in bresnahanOligopoly1982 and matsumuraResolving2023, extends it to more general settings and shows that the condition is not only sufficient but also necessary.
@@
Hereafter, we provide the proof of Theorem 1. The proof of Theorem 1 has the following steps. First, we derive the necessary condition for non-identification based on Definition 1. By rewriting Definition 1, we obtain the transformation that transforms the derivatives of $MC$ into the derivatives of another marginal cost function $MC^{*}$ that leads to observational equivalence given an inverse demand function (Lemma 2). However, the transformation is not valid because it consists of the derivatives of the inverse demand function, that is, it depends on the demand shifter $\bm{X}^{d}$. Then, without any treatment for the demand shifters, $MC^{*}$ depends on the demand shifter $\bm{X}^{d}$, which violates Assumption 1. Therefore, to make the transformation valid, we need to remove the effect of the demand shifter from the transformation, which puts restrictions on the inverse demand function. Then, the restriction leads to a differential equation that characterizes the inverse demand function that leads to non-identification (Proposition 1). Then, to show sufficiency, under the inverse demand function characterized in Proposition 1, we construct a transformation that leads to observational equivalence (Proposition 2).
@@ 3.3 Necessary Condition for the Non-identification
See Appendix A for the detailed proof. The subscript $i$ in $C_{i}$ and $D_{i}$ indicates that $C_{i}$ and $D_{i}$ consist of the derivative of the inverse demand function with respect to $X^{d}_{i}$. Unlike Lemma 1, we have only the transformation relating to marginal cost, and hence we can see the potential obstacle in this transformation more clearly: for the transformations to hold universally for any $Q^{e}$ and $\bm{X}^{s}$, they must be independent of $\bm{X}^{d}$. If the dependence holds, while a demand rotation instrument does not change the equilibrium quantity $Q^{e}$, it changes $MC^{*}$ via $C_{i}$ and $D_{i}$. However, this means that the marginal cost funciton is af funciton of the demand shifter $\bm{X}^{d}$, which is prohibited by Assumption 1. Therefore, in order for the transformations to be valid, both $C_{i}$ and $D_{i}$ must be independent of $\bm{X}^{d}$.
To see when the terms $C_{i}$ and $D_{i}$ independent of $\bm{X}^{d}$, we take the derivative of $C_{i}$ and $D_{i}$ with respect to $X^{d}_{j}$ and put them as zero for all $j=1,\ldots,K^{d}$. A concern is that, as $C_{i}$ and $D_{i}$ already consist of the second-order derivative of the inverse demand function, taking the derivative of $C_{i}$ and $D_{i}$ leads to three-times derivative of the inverse demand function. This is why we need a stronger restriction (Assumption 3) on the inverse demand function than Lau considers (the twice-continuous differentiability assumption in Lau’s claim).
@@
Before we characterize the inverse demand function that makes $C_{i}$ and $D_{i}$ independent of $\bm{X}^{d}$, we first check when $C_{i}$ and $D_{i}$ are not well-defined. This implies that (22) and (24) are violated, and hence the identification holds by Corollary 1. All derivatives in $C_{i}$ and $D_{i}$ are finite by Assumption 3, and hence the numerator and the denominator of $C_{i}$ and $D_{i}$ cannot take a infinite value. In this case, no matter what value of the numerator of $C_{i}$ and $D_{i}$ is, (22) and (24) are violated when at least one of the denominators of $C_{i}$ and $D_{i}$ for $i=1,\ldots,K_{d}$ is zero. In this case, $C_{i}$ and $D_{i}$ will be infinite or indeterminate.
@@
which is the derivative of the marginal revenue under $\theta$ with respect to $X^{d}_{i}$. Therefore, this implies that the marginal revenue under $\theta$ is not affected by the change in $X^{d}_{i}$ at $(\tilde{Q},\tilde{\bm{X}}^{d})$. This also implies that the equilibrium quantity is not affected by the change in $\tilde{X}^{d}_{i}$ at $(\tilde{Q},\tilde{\bm{X}}^{d})$. Thus, $\tilde{X}^{d}_{i}$ has the same role as a demand rotation instrument at $(\tilde{Q},\tilde{\bm{X}}^{d})$ under $\theta$.
The non-identification implies that (25) does not hold for any $Q^{e}$ and $\bm{X}^{d}$. Once we have $(\tilde{Q},\tilde{\bm{X}}^{d})$ for some $i$ where (25) holds, it is enough for the identification. While it is hard to characterize the inverse demand function that satisfies (25) for some $(\tilde{Q},\tilde{\bm{X}}^{d})$ and for some $i$, we can characterize the inverse demand function that satisfies (25) for all $(Q,\bm{X}^{d})$. Note that (25) is a partial differential equation, and we can be solved analytically. Thus, we can characterize the inverse demand function that always leads to identification for any $Q$ and $\bm{X}^{d}$:
@@
Suppose that $\theta\neq 0$. Then, the conduct parameter and the marginal cost function are identified when the inverse demand function is such that
@@
where $\bm{X}^{d}_{-\mathcal{I}}$ is the vector of demand shifters whose index is not in $\mathcal{I}$, and $r:\mathbb{R}^{K_{d}}\rightarrow\mathbb{R}$ and $s:\mathbb{R}^{|\bm{X}^{d}_{-\mathcal{I}}|+1}$ are twice continuously differentiable functions.
@@
Therefore, both imply that when (25) holds, the equilibrium quantity is not affected by the change in $X^{d}_{i}$, and hence $X^{d}_{i}$ always works as a demand rotation instrument under $\theta$. On the other hand, for any other $\theta^{*}\neq\theta$, the equilibrium condition becomes
@@
Now, suppose that (25) does not hold for any $Q$ and $\bm{X}^{d}$ and for all demand shifters. Then, we take the derivative of $C_{i}$ and $D_{i}$ with respect to $X^{d}_{j}$ for all $j=1,\ldots,K_{d}$ and put the derivative being equal to zero. This process leads to differential equations of the inverse demand functions. Fortunately, these differential equations have analytical solutions, and we can characterize an inverse demand function that satisfies the independence of $C_{i}$ and $D_{i}$ of the demand shifter:
Proposition 1.
@@
where $\alpha\neq-\frac{1}{\theta}$ is a constant, and $r:\mathbb{R}\rightarrow\mathbb{R}$ and $s:\mathbb{R}_{+}\rightarrow\mathbb{R}$ are twice continuously differentiable functions.
See Appendix A for the proof. This proposition characterizes the necessary condition of the non-identification. As we have discussed in the introduction, this inverse demand function does not allow for any demand rotation instrument. Recall that bresnahanOligopoly1982 provides the idea of using a demand rotation instrument to identify the conduct parameter, matsumuraResolving2023 formalize his idea by deriving a sufficient condition that guarantees the inverse demand function includes demand rotation instruments for the non-identification in Bresnahan’s setting. The contraposition of their sufficient condition implies that the non-identification holds only if the inverse demand function does not include any demand shifter. As the current environment is more general than bresnahanOligopoly1982 and matsumuraResolving2023, Proposition 2 is the refinement of the result in matsumuraResolving2023 in general settings.
A difference from Lau’s claim is that the conduct parameter can be identified even when the demand shifter is a scalar as long as the inverse demand function includes a demand rotation instrument. For example, consider an inverse demand function with a scalar demand rotation instrument:
@@
This case does not correspond to (15). Note also that $C_{1}$ and $D_{1}$ are given as
@@
This implies that the transformation (20) is not valid for any marginal cost function because $D_{1}(Q,X^{d}_{1})$ depends on $X^{d}_{1}$. Therefore, any $MC^{*}$ that leads to observational equivalence with $MC$ should depend on the demand shifter, under which the identification is possible.
@@ 3.4 Sufficient Condition for Non-identification
We turn to check the sufficiency of the inverse demand function characterized in Proposition 1. To see whether the inverse demand function (15) can lead to the non-identification, we check whether there exists a transformation between the equilibrium conditions under $\mathcal{E}=(\theta,MC)$ and $\mathcal{E}^{*}=(\theta^{*},MC^{*})$. In other words, we check whether there exists a transformation $T$ such that
@@
Suppose that there exists such a transformation $T$ and that an equilibrium quantity $Q^{e}$ satisfies the equilibrium condition under $\mathcal{E}$. Under $\mathcal{E}^{*}$, we have that
@@
The inside of $T$ is the equilibrium condition under $\mathcal{E}$. Therefore, when $T(0)=0$, the above implies that $Q^{e}$ also satisfies the equilibrium condition under $\mathcal{E}^{*}$, and hence the non-identification holds.
In fact, under the inverse demand function (15), we can construct the transformation $T$ by using (20). Therefore, we can have the following sufficient condition for the non-identification:
Proposition 2.
Given the inverse demand function (15) where $\alpha\neq-\frac{1}{\theta}$, there exists a transformation $T$ between the equilibrium conditions under $\mathcal{E}$ and $\mathcal{E}^{*}$. Therefore, non-identification of the conduct parameter and the marginal cost function holds.
The proof is given in Appendix A. While Lau shows the separable inverse demand function is also a sufficient condition of the non-identification, Lau’s sufficiency proof is incomplete because it can be verified that his proof implicitly assumes that there exists a transformation $T$ under a separable inverse demand function (8). However, the separability itself does not guarantee the existence of the transformation. In contrast, Proposition 2 shows that the inverse demand function without any demand rotation instrument can lead to a transformation that leads to non-identification.
@@ 3.5 Discussion
Rethinking the role of demand rotation instrument:
Proposition 1 emphasizes the role of demand rotation instruments. To get the intuition of how the instrument relates to the proof, consider an economy with a linear inverse demand with a demand rotation instrument and a linear marginal cost. bresnahanOligopoly1982 considers this environment. Then, assume that the true conduct is the perfect competition, that is, $\theta=0$, and the alternative conduct is the monopoly, that is, $\theta^{*}=1$. Figure 1 illustrates this environment. Note that the equilibrium quantity under perfect competition is determined by the intersection of the inverse demand function and the marginal cost function $MC$. The equilibrium quantity under the monopoly is determined by the intersection of the marginal revenue $MR^{*}$ and the marginal cost function $MC^{*}$.
@@
In Figure 1(a), the perfect competition and the monopoly lead to the same equilibrium $E$. Then, consider the change in a demand rotation instrument. In Figure 1(b), the demand rotation instrument changes the intercept and slope of the demand function without changing the equilibrium point under the true model. Under the monopoly, the new equilibrium point is $E^{*}$, which is different from $E$, and hence the observational equivalence is violated. So as to lead to the same equilibrium after the change in the demand rotation instrument ($E=E^{*})$, $MC^{*}$ should shift as in Figure 1(c) and 1(d). Note that we changed only the demand shifter, and hence the cost shifter is unchanged. Thus, for the observational equivalence to hold, the shift in $MC^{*}$ should be derived from the change in the demand rotation instrument, which implies that $MC^{*}$ is a function of the demand shifter. However, this is impossible because it violates Assumption 1, that is, the demand shifter should not affect the marginal cost function. The transformation in Lemma 2 clearly shows the possibility of the dependence of the marginal cost function on demand shifters, and hence we need to shut down the effect of demand shifters, which leads to the inverse demand function (15).
@@
where $\mu(Q,\bm{X}^{s})$ depends only on $Q$ and $\bm{X}^{s}$. This corresponds to Equation (15) in lauIdentifying1982. Then he concludes that there is a transformation $T$ such that
@@
The existence of the transformation $T$ is shown by integrating (37) with respect to the cost shifter when it is a scalar, and by applying Lemma B.1 from goldmanNote1964 when the cost shifter is a vector. Note also that in our approach, both (17) and (18) are necessary to derive the transformation between marginal cost functions. In contrast, while Lau derives both equations (17) and (18) in his proof, he relies solely on (18) to show the non-identification.
While Lau does not explicitly define $\mu$ in his formulation, it can be verified that his $\mu$ corresponds to our $\lambda$. However, he neglects a point that the function $\mu$ implicitly depends on the demand shifter $\bm{X}^{d}$, and this dependence is not addressed in his analysis. As long as the effect of the demand shifter cannot be eliminated from $\mu$, the transformation $T$ between $MC$ and $MC^{*}$ will necessarily depend on $\bm{X}^{d}$, which violates Assumption 1. Unfortunately, Lau does not provide a justification for why the effect of the demand shifter can be removed from $\mu$, and hence it leaves a critical gap in the argument. Furthermore, Lau assumes that the inverse demand functions are twice continuously differentiable, which allows him to apply Lemma B.1 to justify the existence of the transformation $T$. To validly eliminate the effect of $\bm{X}^{d}$ from $\mu$, it is necessary to differentiate $\mu$ with respect to $\bm{X}^{d}$ and is required that the derivative is equal to zero. This step, in turn, requires that the inverse demand function be at least three-times continuously differentiable as in our proof. Therefore, the twice-continuous differentiability is not enough to characterize the non-identification.
4 Identification of All Model Components
So far, as in lauIdentifying1982, we have characterized the non-identification problem by assuming that the inverse demand function is already identified and then have not discussed the identification of the inverse demand function given the data. When the inverse demand function is also identified, all components in the conduct parameter model can be identified. Therefore, we check whether the inverse demand function, where the conduct parameter and the marginal cost function are not identified, can be identified. Unfortunately, we have the following proposition:
Proposition 3.
An inverse demand function cannot be identified when it has the following form:
$P(Q,\bm{X}^{d})=Q^{\alpha}r(\bm{X}^{d})+s(Q),$ (39)
where $\alpha$ is a constant, and $r$ and $s$ are twice continuously differentiable functions.
[…]
Another unfortunate fact is that demand rotation instruments do not help identify an inverse demand function. The following example demonstrates this point, and hence demand rotation instruments cannot be a sufficient condition for the identification of the inverse demand function:
Example 1.
Suppose the inverse demand function is given as
$P(Q,\bm{X}^{d})=Q^{\alpha}r(\bm{X}^{d})+s(Q)+r(\bm{X}^{d}).$ (40)
Here, $\bm{X}^{d}$ can work as a demand rotation instrument through $r$. However, in the same way as in Proposition 3, it is easy to see that for any $r$ and $s$, the inverse demand functions with $\tilde{r}(\bm{X}^{d})\equiv r(\bm{X}^{d})+1$ and $\tilde{s}(Q)\equiv s(Q)-Q^{\alpha}-1$ lead to the same price for any quantity and demand shifter, which implies that the inverse demand function is not identified.
Therefore, our results so far imply that the identification of the conduct parameter is guaranteed once the inverse demand function includes demand rotation instrument, but the inverse demand function with demand rotation instruments should be identified in advance.

### v3

@@ 3.2 Main Result
We now show the following theorem:
@@
Given Assumption 1, 2, 3, and 4, and data on price, quantity, and other exogenous variables,
  • (i)
    if the conduct parameter $\theta$ and the marginal cost function $MC$ are not identified, the industry inverse demand function must be given as
@@
    where $\alpha\neq-\frac{1}{\theta}$.
  • (ii)
    if the inverse demand function takes the form (13) with the properties described in (i), then the conduct parameter and the marginal cost are not identified.
@@
As in Lau’s claim, when the demand shifter is a vector, the inverse demand function is separable because the demand shifters affect the inverse demand function only through $r$. However, our result tells more about under what type of separable function the non-identification holds and has a clear economic interpretation. As in Claim 1, our result has a special case where any change in demand shifters does not affect the marginal revenue under the true conduct but does under any false conduct, and hence the identification always holds ($\alpha=-\frac{1}{\theta}$). Except in this case, observe that the demand shifter changes only the slope of the inverse demand function when $\alpha\neq 0$ or changes only the intercept of the inverse demand function when $\alpha=0$. Recall that demand rotation instruments can change the intercept and the slope of the inverse demand function simultaneously. Therefore, we cannot have any demand rotation instruments under (13). The result implies that the identification holds if and only if the inverse demand function has demand shifters that work as demand rotation instruments. Here, we are explicit that a demand rotation instrument is not necessarily a single variable, but could be a combination of changes in some demand shifters that alter the slope and the intercept of the inverse demand function simultaneously. Note also that Bresnahan graphically demonstrates that the conduct parameter is identified when a demand rotation instrument moves the demand curve in a way that leaves the equilibrium point unchanged at the true conduct but changes the equilibrium under any false conduct. However, this is an extreme case and it suffices that the instrument changes the slope and the intercept simultaneously. This clearly reinforces the idea of Bresnahan (1982) using demand rotation instruments to identify the conduct parameter in more general settings.
@@
Hereafter, we provide the proof of Theorem 1. First, we derive the necessary condition for non-identification based on Definition 1. By rewriting Definition 1, we obtain a transformation that maps the derivatives of $MC$ into the derivatives of another marginal cost function $MC^{*}$ that yields observational equivalence given an inverse demand function (Lemma 2). However, the transformation is not valid because it consists of derivatives of the inverse demand function and hence depends on the demand shifter $\bm{X}^{d}$. Without additional restrictions, $MC^{*}$ would therefore depend on $\bm{X}^{d}$, violating Assumption 1. Therefore, to make the transformation valid, we need to remove the effect of the demand shifter from the transformation, which puts restrictions on the inverse demand function. Then, the restriction leads to a differential equation that characterizes the inverse demand function that leads to non-identification (Lemma 4). Then, to show sufficiency, under the inverse demand function characterized in Lemma 4, we construct a transformation that leads to observational equivalence (Lemma 5).
@@ 3.3 Necessary Condition for the Non-identification
See Appendix A for the detailed proof. The subscript $i$ in $C_{i}$ and $D_{i}$ indicates that $C_{i}$ and $D_{i}$ consist of the derivative of the inverse demand function with respect to $X^{d}_{i}$. Unlike Lemma 1, we have only the transformation relating to marginal cost, and hence we can see the potential obstacle in this transformation more clearly: for the transformations to hold universally for any $Q^{e}$ and $\bm{X}^{s}$, they must be independent of $\bm{X}^{d}$. If the dependence holds, while a demand rotation instrument does not change the equilibrium quantity $Q^{e}$, it changes $MC^{*}$ via $C_{i}$ and $D_{i}$. However, this means that the marginal cost function is a function of the demand shifter $\bm{X}^{d}$, which is prohibited by Assumption 1. Therefore, in order for the transformations to be valid, both $C_{i}$ and $D_{i}$ must be independent of $\bm{X}^{d}$.
To see when the terms $C_{i}$ and $D_{i}$ are independent of $\bm{X}^{d}$, we take the derivative of $C_{i}$ and $D_{i}$ with respect to $X^{d}_{j}$ and set them equal to zero for all $j=1,\ldots,K_{d}$. A concern is that, as $C_{i}$ and $D_{i}$ already consist of the second-order derivative of the inverse demand function, taking the derivative of $C_{i}$ and $D_{i}$ leads to three-times derivative of the inverse demand function. This is why we need a stronger restriction (Assumption 3) on the inverse demand function than Lau considers (the twice-continuous differentiability assumption in Lau’s claim).
@@
Before we characterize the inverse demand function that makes $C_{i}$ and $D_{i}$ independent of $\bm{X}^{d}$, we first check when $C_{i}$ and $D_{i}$ are not well-defined. This implies that (20) and (22) are violated, and hence the identification holds by Corollary 1. All derivatives in $C_{i}$ and $D_{i}$ are finite by Assumption 3, and hence the numerator and the denominator of $C_{i}$ and $D_{i}$ cannot take an infinite value. In this case, regardless of the numerator of $C_{i}$ or $D_{i}$, (20) and (22) are violated whenever at least one denominator for $i=1,\ldots,K_{d}$ is zero. In this case, $C_{i}$ and $D_{i}$ will be infinite or indeterminate.
@@
which is the derivative of the marginal revenue under $\theta$ with respect to $X^{d}_{i}$. Therefore, this implies that the marginal revenue under $\theta$ is not affected by the change in $X^{d}_{i}$ at $(\tilde{Q},\tilde{\bm{X}}^{d})$. This also implies that the equilibrium quantity is not affected by the change in $\tilde{X}^{d}_{i}$ at $(\tilde{Q},\tilde{\bm{X}}^{d})$.
The non-identification implies that (23) does not hold for any $Q^{e}$ and $\bm{X}^{d}$. Once we have $(\tilde{Q},\tilde{\bm{X}}^{d})$ for some $i$ where (23) holds, it suffices for identification. While it is hard to characterize the inverse demand function that satisfies (23) for some $(\tilde{Q},\tilde{\bm{X}}^{d})$ and for some $i$, we can characterize the inverse demand function that satisfies (23) for all $(Q,\bm{X}^{d})$. Note that (23) is a partial differential equation that can be solved analytically. Thus, we can characterize the inverse demand function that always leads to identification for any $Q$ and $\bm{X}^{d}$:
@@
Suppose that $\theta\neq 0$ and $Q>0$. Then, the conduct parameter and the marginal cost function are identified when the inverse demand function is such that
@@
where $\bm{X}^{d}_{-\mathcal{I}}$ is the vector of demand shifters whose index is not in $\mathcal{I}$, and $r:\mathbb{R}^{K_{d}}\rightarrow\mathbb{R}$ and $s:\mathbb{R}^{|\bm{X}^{d}_{-\mathcal{I}}|+1}$ are at least twice-continuously differentiable.
@@
Therefore, when (23) holds, the equilibrium quantity is not affected by the change in $X^{d}_{i}$. On the other hand, for any other $\theta^{*}\neq\theta$, the equilibrium condition becomes
@@
Now, suppose that (23) does not hold for any $Q$ and $\bm{X}^{d}$ and for all demand shifters. Then, we take the derivative of $C_{i}$ and $D_{i}$ with respect to $X^{d}_{j}$ for all $j=1,\ldots,K_{d}$ and put the derivative being equal to zero. This process leads to differential equations for the inverse demand function. Fortunately, these differential equations have analytical solutions, and we can characterize an inverse demand function that satisfies the independence of $C_{i}$ and $D_{i}$ of the demand shifter:
Lemma 4.
@@
where $\alpha\neq-\frac{1}{\theta}$, and $r:\mathbb{R}^{K_{d}}\rightarrow\mathbb{R}$ and $s:\mathbb{R}_{+}\rightarrow\mathbb{R}$ are at least twice-continuously differentiable.
See Appendix A for the proof. Technically, the twice-continuously differentiability of $r$ and $s$ is required to meet Assumption 3. As we have discussed, this inverse demand function does not allow for any demand rotation instrument. Recall that Bresnahan (1982) provides the idea of using a demand rotation instrument to identify the conduct parameter. Then, Matsumura and Otani (2023) formalize his idea by deriving a sufficient condition that guarantees the inverse demand function includes demand rotation instruments for the identification in Bresnahan’s setting. The contraposition of their sufficient condition implies that the non-identification holds only if the inverse demand function does not include any demand rotation instrument. Therefore, our characterization is the generalization of their result in general settings.
Another difference from Lau’s claim is that the conduct parameter can be identified even when the demand shifter is a scalar as long as the inverse demand function includes a demand rotation instrument. For example, consider an inverse demand function with a scalar demand rotation instrument:
@@
In this case, $C_{1}$ and $D_{1}$ are given as
@@
This implies that the transformation (18) is not valid for any marginal cost function because $D_{1}(Q,X^{d}_{1})$ depends on $X^{d}_{1}$. Therefore, any $MC^{*}$ that leads to observational equivalence with $MC$ should depend on the demand shifter, under which the exclusion restriction is violated and hence identification is possible.
Remark 2.
The differentiability assumptions are imposed on the inverse demand function through Assumption 3. The functions $r$ and $s$ appear as components of the solution to a differential equation implied by the non-identification condition. Their differentiability is therefore inherited from that of $P$ and is not imposed as an independent assumption. Cases in which $r$ or $s$ fail to be sufficiently differentiable are outside the scope of the theorem.
@@ 3.4 Sufficient Condition for Non-identification
We turn to check the sufficiency of the inverse demand function characterized in Lemma 4. To see whether the inverse demand function (13) can lead to non-identification, we check whether there exists a transformation between $MC$ and $MC^{*}$ and the marginal revenue under $\theta$ and under $\theta^{*}$ such that any equilibrium in $\mathcal{E}$ can also be an equilibrium in $\mathcal{E}^{*}$ by using the mapping. Formally, suppose that there exists a transformation $T$ such that
@@
Then, suppose also that an equilibrium quantity $Q^{e}$ satisfies the equilibrium condition under $\mathcal{E}$. When we have that for $\mathcal{E}^{*}$,
@@
then $Q^{e}$ also satisfies the equilibrium condition under $\mathcal{E}^{*}$, and hence the non-identification holds. In fact, under the inverse demand function (13), we can construct such transformation $T$ by using (18). Therefore, we can have the following sufficient condition for the non-identification:
Lemma 5.
Given the inverse demand function (13) where $\alpha\neq-\frac{1}{\theta}$, given a conduct parameter and a marginal cost function, we can construct another pair of a conduct parameter and a marginal cost function that leads to the same equilibrium point for any value of demand shifters and cost shifters. Therefore, non-identification of the conduct parameter and the marginal cost function holds.
The proof is given in Appendix A. While Lau shows the separable inverse demand function is also a sufficient condition for non-identification, his sufficiency proof is incomplete because it can be verified that his proof implicitly assumes that there exists a transformation $T$ under a separable inverse demand function (6). However, the separability itself does not guarantee the existence of the transformation. In contrast, Lemma 5 shows that the inverse demand function without any demand rotation instrument can lead to a transformation that leads to non-identification.
@@ 3.5 Discussion
Rethinking the role of demand rotation instruments:
Lemma 4 emphasizes the role of demand rotation instruments. Intuitively, when a change in a demand shifter keeps the equilibrium quantity the same, we have two equilibrium conditions:
$P(Q,\bm{X}^{d})+\theta Q\frac{\partial P}{\partial Q}(Q,\bm{X}^{d})$ $=MC(Q,\bm{X}^{s})$ (35)
$P(Q,\tilde{\bm{X}}^{d})+\theta Q\frac{\partial P}{\partial Q}(Q,\tilde{\bm{X}}^{d})$ $=MC(Q,\bm{X}^{s})$ (36)
Here, we have two equations and two unknowns $(\theta,MC(Q,\bm{X}^{s}))$, and hence we can solve the system of equation with respect to the conduct parameter and the marginal cost function. Therefore, the existence of a demand shifter that keep the equilibrium quantity the same is confirmed as the sufficient condition for the identification in more general setting.
To build intuition for how the instrument relates to the proof, consider an economy with a linear inverse demand with a demand rotation instrument and a linear marginal cost. Bresnahan (1982) considers this environment. Then, assume that the true conduct is the perfect competition, that is, $\theta=0$, and the alternative conduct is the monopoly, that is, $\theta^{*}=1$. Figure 1 illustrates this environment. Note that the equilibrium quantity under perfect competition is determined by the intersection of the inverse demand function and the marginal cost function $MC$. The equilibrium quantity under the monopoly is determined by the intersection of the marginal revenue $MR^{*}$ and the marginal cost function $MC^{*}$.
@@
In Figure 1(a), the perfect competition and the monopoly lead to the same equilibrium $E$. Then, consider the change in a demand rotation instrument. In Figure 1(b), the demand rotation instrument changes the intercept and slope of the demand function without changing the equilibrium point under the true model. Under the monopoly, the new equilibrium point is $E^{*}$, which is different from $E$, and hence the observational equivalence is violated. To obtain the same equilibrium after the change in the demand rotation instrument ($E=E^{*}$), $MC^{*}$ should shift as in Figures 1(c) and 1(d). Note that we changed only the demand shifter, and hence the cost shifter is unchanged. Thus, for the observational equivalence to hold, the shift in $MC^{*}$ should be derived from the change in the demand rotation instrument, which implies that $MC^{*}$ is a function of the demand shifter. However, this is impossible because it violates Assumption 1, that is, the demand shifter should not affect the marginal cost function. The transformation in Lemma 2 clearly shows the possibility of the dependence of the marginal cost function on demand shifters, and hence we need to shut down the effect of demand shifters, which leads to the inverse demand function (13).
@@
where $\mu(Q,\bm{X}^{s})$ depends only on $Q$ and $\bm{X}^{s}$. This corresponds to Equation (15) in Lau (1982). Then he concludes that there is a transformation $T$ such that
@@
The existence of the transformation $T$ is shown by integrating (37) with respect to the cost shifter when it is a scalar, and by applying Lemma B.1 from Goldman and Uzawa (1964) when the cost shifter is a vector. Note also that in our approach, both (15) and (16) are necessary to derive the transformation between marginal cost functions. In contrast, while Lau derives both equations (15) and (16) in his proof, he relies solely on (16) to show the non-identification.
While Lau does not explicitly define $\mu$ in his formulation, it can be verified that his $\mu$ corresponds to our $\lambda$. However, he neglects a point that the function $\mu$ depends on the demand shifter $\bm{X}^{d}$, and this dependence is not addressed in his analysis. As long as the effect of the demand shifter cannot be eliminated from $\mu$, the transformation $T$ between $MC$ and $MC^{*}$ could depend on $\bm{X}^{d}$, which violates Assumption 1. Additionally, separability does not resolve the dependence of $\mu$ on the demand shifter. Unfortunately, Lau does not provide a justification for why he can assume that the effect of the demand shifter can be removed from $\mu$, and hence it leaves a critical gap in the argument. Additionally, Lau only assumes that the inverse demand functions are twice continuously differentiable, which allows him to apply Lemma B.1 to justify the existence of the transformation $T$. However, to eliminate the effect of $\bm{X}^{d}$ from $\mu$, it is necessary to differentiate $\mu$ with respect to $\bm{X}^{d}$ and require that the derivative is equal to zero as in our analysis. This step, in turn, requires that the inverse demand function be at least three-times continuously differentiable as in our proof. Therefore, twice-continuous differentiability is not enough to characterize the non-identification.